### v1

@@ Abstract
This paper quantifies the causal effect of cohesion policy on EU regional output and investment focusing on one of its least studied instruments, i.e., the Cohesion Fund (CF). We employ modern causal inference methods to estimate not only the local average treatment effect but also its time-varying and heterogeneous effects across regions. Utilizing this method, we propose a novel framework for evaluating the effectiveness of CF as an EU cohesion policy tool. Specifically, we estimate the time varying distribution of the CF’s causal effects across EU regions and derive key distribution metrics useful for policy evaluation. Our analysis shows that relying solely on average treatment effects masks significant heterogeneity and can lead to misleading conclusions about the effectiveness of the EU’s cohesion policy. We find that the impact of the CF is frontloaded, peaking within the first seven years after a region’s initial inclusion in the program. The distribution of the effects during this first seven-year cycle of funding is right skewed with relatively thick tails. This indicates positive effects but unevenly distributed across regions. Moreover, the magnitude of the CF effect is inversely related to a region’s relative position in the initial distribution of output, i.e., relatively poorer recipient regions experience higher effects compared to relatively richer regions. Finally, we find a non-linear relationship with diminishing returns, whereby the impact of CF declines as the ratio of CF funds received to a region’s gross value added (GVA) increases.
@@ 1 Introduction
Cohesion policy is one of the most important fiscal policy initiatives at the European Union (EU) level. Over the last three decades, nearly €960 billion (in 2015 prices) have been disbursed to the EU regions as part of this sub-national transfers program. To put things in a historical perspective, cohesion-related expenditures account for almost one-third of the EU’s budget (see Benedetto (2024) and Bachtler (2022)), making it one of the largest redistribution programs ever implemented in the EU. In the current seven-year budget period, [^1] The so-called Multiannual Financial Framework. €392 billion have been dedicated to cohesion policy objectives, reinforcing its status as a top spending priority of the EU. Naturally, given its scale and overarching objectives, the effectiveness of cohesion policy has been debated over time by both academics and policymakers (see European Commission (2024)).
@@
This paper quantifies the causal effect of cohesion policy on EU regional output and investment. By adopting recent developments in the causal inference and potential outcomes literature, i.e., Xu (2017), Athey et al. (2021) and Borusyak et al. (2024), [^2] For a recent review see Xu (2023) we extend beyond the existing literature (e.g., Becker et al. (2010, 2013, 2018)) which has primarily focused on estimating a common effect across all treated regions, i.e. the (local) average treatment effect. Our empirical strategy allows us to capture not only the average impact of cohesion policy but also its time-varying and heterogeneous effects across regions, providing a more nuanced understanding of its effectiveness.
We focus on one of the least studied instruments of cohesion policy: the Cohesion Fund (CF). This policy instrument is solely available to the relatively poorer Member States and has disbursed, over time, almost 20% of total cohesion policy expenditures. Our paper introduces a novel approach to evaluate its core objectives, which go beyond simply boosting income across recipient regions to actively address economic disparities among EU regions. To do this, we utilize the strengths of our methodology, which allows us to estimate the impact of each region for every year following its inclusion in the CF program. This enables us to estimate the time-varying distribution of CF effects across regions and derive key distributional metrics that are useful for policy evaluation. Specifically, we compute the percentage of regions that experience a positive (or negative) impact. Moreover, we examine the skewness of the distribution which indicates whether the CF program has predominantly positive or negative effects, along with the kurtosis which measures the degree of uneven dispersion of CF effects across regions. We do this over consecutive funding cycles aligning with the EU’s institutional seven-year budgeting framework. In the context of the CF’s policy evaluation, a decline in the percentage of regions experiencing a positive impact over time may suggest a diminishing effectiveness of the CF. However, this should not always be interpreted at face value but, rather, should be assessed alongside additional distributional metrics. Specifically, a simultaneous decrease in skewness and/or kurtosis—indicating a more symmetric distribution—may instead suggest that the CF effect becomes more symmetric as a result of reducing regional disparities over time.
The identification of the causal effects of such regional transfer programs has mainly been conducted by relying either on panel data methods related to the Differences-in-Differences (DiD) approach (see for example Gobillon and Magnac (2016) for a discussion) or by employing a Regression Discontinuity Design (RDD) to assess, for example, the impact of the cohesion policy expenditures across the regions of the EU. These papers leverage on the institutional rule that determines whether a region is eligible or not for the relevant EU objectives (e.g. the so-called Objective I or Convergence Objective), which in turn determines the corresponding size of the cohesion transfers - e.g. see Becker et al. (2010, 2012) - more details on this are provided in section 2.
However, in the analysis of regional policies there are specific issues which require careful consideration. The first issue, which has received quite a lot of attention, is that of the spatial dependence across units. These dependencies arise mainly due to geographical reasons and are local in nature. A number of studies have employed standard spatial econometrics techniques to account for these issues (see, among others, Amendolagine et al. (2024), Bourdin (2018), Crescenzi and Giua (2020), Di Caro and Fratesi (2022) and Mohl and Hagen (2010)) to account for these factors. To our knowledge, only Crescenzi and Giua (2020) have attempted to account for these dependencies in a causal inference contaxt, by employing a spatial RDD in order to estimate the causal effects of cohesion policy.
A second important issue is that of cross-sectional dependence, highlighted by Gobillon and Magnac (2016). Specifically, cross-sectional dependence is a form of time-varying heterogeneity which arises from the presence of unobserved common factors whose impact extends well beyond a local geographical area. Examples of such factors that are prominent in our sample include the Global Financial Crisis of 2008 and the debt crisis of the southern periphery of the EU. These unobserved common factors can lead to severe biases if they are correlated with the treatment variable.
Lastly, it is important to recognize that regional policies are more likely than not to exert region-specific effects. To our knowledge, the relevant empirical literature on the policy evaluation of Cohesion transfers has largely assumed that the impact of the EU’s regional policy is identical across all regions. Essentially, this amounts to disregarding any heterogeneity that arises due to differences in the relative position of regions across the income distribution along with a wide range of other idiosyncratic characteristics.
In this paper, we overcome these challenges by drawing on recent developments from the causal inference and potential outcomes literature in order to estimate the causal effects of the Cohesion Fund payments. Specifically, our analysis builds on the matrix completion technique implemented in the context of factor models - see, Athey et al. (2021) and Xu (2017). The factor or interactive fixed effects models (see Bai (2009a), Pesaran (2006)) nest the standard Two-Way Fixed Effects model and can directly account for the presence of cross-sectional dependence and the potential correlation between the treatment and the outcome (see Gobillon and Magnac (2016)).
Our approach is a generalization of the synthetic control method of Abadie et al. (2010), which relaxes key assumptions of standard causal inference methods. In particular, by employing the synthetic control methodology we avoid the imposition of the so-called parallel trends assumption which is central to all the techniques that rely on the Differences-in-Differences approach -see, Angrist and Pischke (2009), Sun and Abraham (2021), De Chaisemartin and d’Haultfoeuille (2024). It is worth mentioning that the presence of the time-varying unobserved common factors, which are an integral part of our approach, would invalidate this assumption (see Xu (2017)). In our setup, the interactive fixed effects will directly model and account for these confounders. Moreover, under this generalized approach, the staggered adoption of the treatment from more than one units is allowed.
Our main findings are as follows. We find that, on average, the CF has a persistent positive effect on both output (Gross Value Added) and investment (Gross Fixed Capital Formation) of the recipient regions, compared to a counterfactual scenario where no funds are provided by the EU. The impact of the CF is frontloaded, peaking within the first seven years after a region’s initial inclusion in the program. For a hypothetical average EU region, the CF led to an increase in GVA per capita of €468 compared to the counterfactual scenario. In the second seven-year cycle, i.e. after 14 years of inclusion in the CF program, the positive effect persists but diminishes in magnitude, with GVA per capita increasing by €330 compared to the counterfactual. In other words, a citizen in this hypothetical EU region will be wealthier by €468 in the 7th year after the inclusion to the program and by €330 in the 14th year. Similarly, the hypothetical average EU region, experiences an increase in GFCF per capita of €385 and €415 after 7 and 14 years, respectively, compared to the counterfactual.
However, our analysis shows that relying solely on average treatment effects masks significant heterogeneity and can lead to misleading conclusions about the effectiveness of this EU policy tool. We provide novel evidence on the distribution of the impact of the Cohesion Funds across EU regions and its evolution over consecutive seven-year cycles (mimicking the EU budgeting cycles). Specifically, the distribution of the GVA effect during the first seven-year cycle of funding is right skewed with relatively thick tails. This indicates positive GVA effects which are unevenly distributed across regions. In the second cycle (after 14 years of treatment) both the skewness and the kurtosis of the distribution decrease. This indicates that the magnitude of the CF effect becomes smaller, while its impact is more symmetric across regions. We quantify the distributional effects and, specifically, we find that in the first seven-year cycle 64.5% of treated regions experience a positive effect, while this percentage reduces to 49% in the second seven-year cycle. When we categorize regions into percentiles based on their initial GVA, we find that the poorest regions – those in the bottom 0-10th percentile– experience a significantly larger in magnitude positive effect from receiving Cohesion Funds. Specifically, their GVA after the first seven-years of receiving the fund is estimated to grow by 21.5 percentage points more compared to the counterfactual scenario. For regions in the 10th-25th and 50th-75th percentile, the gains are 18.8 and 12.3 percentage points higher, respectively, compared to the counterfactual. However, the benefits diminish for the relatively wealthier regions and even turn to negative for the wealthiest EU regions. Another important finding is that the positive effect persists in the second seven-year cycle, i.e., after 14 years under treatment; however, it becomes significantly smaller in magnitude. Turning to the GFCF distribution, most of the patterns observed for GVA hold. However, we observe two key differences. First, a higher percentage of the treated regions experience a positive effect within the first seven years under treatment, namely 90%. Second, the distribution is relatively more right-skewed and leptokurtic (higher tails) which implies that the positive effects are larger and more unevenly distributed across region. For the poorest regions, i.e., those in the 0-10th percentile in terms of initial GVA per capita, the investment effect is 57.5 percentage points higher compared to the counterfactual. This impact is more than 2.5 larger than the estimated effect on GVA growth. This finding provides evidence that the CF payments have a regional impact that adheres to the policy objectives, namely, to stimulate regional investment and infrastructure, and, through these channels, enhance regional GVA growth, especially for the poorest regions in the EU.
Having these granular effects of the Cohesion Fund payments by region and over time, we use them to visualize the effects on a map of the European regions. A distinct spatial pattern emerges from visual inspection of such a map. The smallest effects of the Cohesion Fund are concentrated in Southern Europe. In contrast, the largest effects are observed in Eastern Europe (note that countries in Northern Europe were not eligible for CF support). This pattern reveals a new geographical division within Europe. Specifically, Southern Europe, which consists of relatively wealthier regions, experiences weaker CF effects, whereas Eastern Europe, which consists of poorer regions that became members of the EU in the early 2000s, benefits the most from such a policy. Moreover, when we aggregate these results at the country level, most of the findings observed at the regional level are confirmed. Specifically, the impact of the Cohesion Fund is frontloaded, with the largest effects estimated within the first seven-year cycle. Additionally, the magnitude of the impact is negatively correlated with a country’s initial GVA per capita, meaning poorer countries tend to experience stronger positive effects. These differences in GVA growth stem from the causal impact of regions within each country receiving Cohesion Fund support. Finally, we find a nonlinear relationship and diminishing returns for the CF impact on GVA as a function of the CF to GVA ratio. That is, for relatively low funding as a share of GVA, the effect is negligible or even negative. As funding increases, the effect grows, peaking when the cohesion fund reaches 0.58% and 0.94% of the region’s GVA for the seven- and fourteen-year period respectively. Beyond a point, we observe diminishing returns, where the effect of the Cohesion Fund gradually declines and eventually becomes zero or even negative at higher funding as a share of GVA.
@@ 2 Related literature
The present paper contributes to the empirical literature studying the effects of the EU’s cohesion policy expenditures on regional economic performance. Based on the approach utilized, this literature can be split into two broad categories: one that employs causal inference techniques and one that focuses on more traditional (spatial) panel econometrics approaches. A schematic review is presented in Table 1.
Starting with the first branch of the relevant literature, the focus is on disentangling the causal impact that cohesion transfers exert on the economic performance of the recipient EU regions. These papers leverage on an institutional rule that determines whether a region is eligible or not for EU transfers. In particular, the rule stipulates that if a region’s GDP per capita (in PPP terms) is less than 75% of the EU average prior to the commencement of the programming period, then it is awarded the ‘Objective I’ or ‘Convergence criterion’ or ‘less-developed’ region status and is thus eligible for the EU transfers. [^3] It should be clarified here that all EU NUTS2 regions are eligible for cohesion policy transfers. However, the bulk of the transfers is directed to the ‘less developed’ regions mentioned above; that is, it is the intensity of the transfers that changes across regions.
By focusing on a binary treatment indicator based on the GDP per capita rule, the literature examined whether being awarded the ‘less-developed’ status had a causal impact on economic performance. In this context, Becker et al. (2010, 2013, 2018) and Crescenzi and Giua (2020) focus on an interesting observation and develop a ‘fuzzy’ RDD in order to assess the impact of cohesion transfers. In particular, they observe that there are cases of non-compliance with the eligibility rule which led to a distinction between a region being eligible for and a recipient of EU transfers; that is, there exist cases were regions were awarded the Objective 1 status despite having a larger level of GDP per capita as well as cases were the regions were not awarded the ‘less-developed’ status despite meeting the criteria. [^4] Interestingly, Pellegrini et al. (2013) use a ‘sharp’ RDD approach. Becker et al. (2010) show that cohesion policy transfers to the lagging regions increase real per capita GDP growth by 1.6% within the same programming period. This translates to a multiplier of 1.2, implying that the policy yields net benefits. Becker et al. (2013) develop a heterogeneous variant of the RDD in order to examine if absorptive capacity –proxied by an index of the quality of government and human capital– is the driving force behind heterogeneous treatment effects. They conclude that heterogeneity in absorptive capacity is crucial, with only one-third of the regions in their sample exhibiting a high enough level that allows them to attain large, positive economic effects. Lastly, Becker et al. (2018) use an updated dataset that allows them to account for the impact of the financial crisis of 2008. They document a weaker impact of the EU’s cohesion policy transfers during the programming period containing the crisis, pointing to a downward pressure that was exerted by the cyclical effects. Additionally, they document that the effects of the policy seem to be contained within one programming period.
@@
A separate branch of this literature shifted the focus from the broad question of whether cohesion transfers have a causal impact on growth to the more granular issue of whether the level of the transfer intensity that regions receive is what matters for economic performance. Technically, these papers exploit the variation in the intensity of the treatment. In this case, the question at hand is whether higher levels of the treatment lead to a higher or lower impact on the outcome compared to lower levels of the treatment. In this context, Hagen and Mohl (2008), Becker et al. (2012) and Cerqua and Pellegrini (2018) employ a generalized propensity score approach following Hirano and Imbens (2004) and estimate the so-called dose-response function. This allows the examination of whether there exists either a minimum necessary level or a maximum desirable level of regional transfer intensity necessary to achieve positive effects from cohesion policy expenditures (e.g. see Becker et al. (2012)). This has important policy implications, as the existence of either bound implies that a redistribution of the EU’s transfer budget could be undertaken, leading to positive effects for more regions. In particular, Becker et al. (2012) identify that the maximum desirable level of transfer intensity for their sample is equal to 1.3% of GDP (and, specifically, of the beginning-of-period GDP). Moreover, they calculate that the optimal transfer intensity –the level of transfers beyond which one additional euro of transfers leads to the generation of less than one euro of additional GDP– is equal to 0.4% of GDP. Similar results are obtained by Cerqua and Pellegrini (2018). On the contrary, Hagen and Mohl (2008) show that for a sample focusing on earlier programming periods, Cohesion transfers have a positive yet statistically insignificant effect thus rendering void the discussion of an ‘optimal dose’ of EU transfers. Overall, the literature utilizing counterfactual policy evaluation techniques seems to provide evidence pointing to a clear positive effect of cohesion policy funds on regional economic growth and employment.
Our work extends this literature by applying a recently developed technique that is able to disentangle region- and time-specific effects for the EU regions that were recipients of the Cohesion Fund payments. This approach allows us to assess the causal effects of the Cohesion Fund under a more general, heterogeneous setting compared to what was implemented in the relevant literature.
The second broad category of the relevant empirical studies focuses on the use of panel data techniques for the assessment of the impact of cohesion policy expenditures on economic performance. As the approach followed in these articles goes beyond the scope of the present paper, we only briefly discuss it here. Interested readers are referred to the papers included in the upper panel of Table 1 for an entry point to this literature. This rather extensive literature employs various panel econometrics techniques and aims to disentangle the (long-run) effects of these sub-national transfers on regional economic performance using mainly instrumental variables and spatial approaches (see, for example, Di Caro and Fratesi (2022) and Mohl and Hagen (2010)). Recent contributions have incorporated into the analysis institutional factors - e.g. see Rodríguez-Pose and Garcilazo (2015). Lastly, a rather recent set of contributions focuses on the estimation of the multiplier effects of Cohesion expenditures - e.g. see Canova and Pappa (2024).
Table 1: Schematic Literature Review
Panel data
Policy evaluation for EU funding impact assessment
Becker et al. (2013) 186-251 NUTS2(p) 1989-2006 Binary GDP pc growth Fuzzy RDD Homogeneous
Becker et al. (2018) 1989-2013 259 NUTS2 Binary GDP pc growth Fuzzy RDD Homogeneous
Crescenzi and Giua (2020) 2000-2014 Binary GDP pc growth Spatial RDD partially Heterogeneous
Hagen and Mohl (2008) 122 NUTS1/NUTS2 1995-2005 Payments GDP pc growth Dose-response (GPS) Homogeneous
  • Homogeneity: Partial means that the paper includes results based on groupings of regions.
  • Weighting matrix: In order to preserve the readability of the table, the various weighting schemes employed are mentioned here: k-nearest neighbor, Gaussian distance decay function,

### v2

@@ Abstract
This paper quantifies the causal effect of cohesion policy on regional output and investment across the EU regions, focusing on one of its least studied instruments, namely, the Cohesion Fund (CF). The novelty of the paper lies in deviating from the standard methodologies employed in the existing literature such as regression discontinuity design (RDD) or instrumental variables (IV). Instead, we adopt a recently developed causal inference approach based on matrix completion, implemented within a factor model framework. This allows us to propose a novel framework for evaluating the effectiveness of the CF as a key EU cohesion policy instrument. Specifically, we estimate the time varying distribution of the CF’s causal effects across EU regions and derive key distribution metrics useful for policy evaluation. Our analysis shows that relying solely on average treatment effects masks significant heterogeneity and can lead to misleading conclusions about the effectiveness of the EU’s cohesion policy. We find that the impact of the CF is frontloaded, peaking within the first seven years following a region’s initial inclusion in the program. The distribution of the effects during this first seven-year cycle of funding is right-skewed with relatively thick tails. This indicates generally positive effects but unevenly distributed across regions. Moreover, the magnitude of the CF’s effect is inversely related to a region’s relative position in the initial regional GVA distribution, i.e., relatively poorer recipient regions experience higher effects compared to relatively richer regions. We find a non-linear relationship with diminishing returns, whereby the impact of the CF begins to decline beyond a certain threshold as the ratio of CF funds received to a region’s gross value added (GVA) increases.
@@ 1 Introduction
Cohesion policy is one of the most important fiscal policy initiatives at the European Union (EU) level. Over the last three decades, nearly €960 billion (in 2015 prices) have been disbursed to the EU regions as part of this sub-national transfers program. To put things in a historical perspective, cohesion-related expenditures account for almost one-third of the EU’s budget (see Bachtler (2022) and Benedetto (2024)), making it one of the largest redistribution programs ever implemented in the EU. In the current seven-year budget period, [^1] The so-called Multiannual Financial Framework - see the Council Regulation 2020/2093 for additional details on the expenditure breakdown for the 2021-2027 budget. €392 billion have been dedicated to cohesion policy objectives, reinforcing its status as a top spending priority of the EU. Naturally, given its scale and overarching objectives, the effectiveness of cohesion policy has been debated over time by both academics and policymakers (see European Commission (2024)).
@@
This paper quantifies the causal effect of EU cohesion policy on regional output (Gross Value Added) and investment (Gross Fixed Capital Formation). We focus on one of the least studied instruments of cohesion policy: the Cohesion Fund (CF). This policy instrument is available only to the relatively poorer Member States and has accounted for nearly 20% of the total cohesion policy expenditures over time. The first novelty of the paper is the adoption of recent developments in causal inference, i.e., Xu (2017), Athey et al. (2021) and Borusyak et al. (2024) to study this type of questions. [^2] For a recent review see Xu (2023) Thus, we extend beyond the existing literature, see for example Becker et al. (2010), Becker et al. (2013), Gagliardi and Percoco (2017) and Crescenzi and Giua (2020) which has primarily focused on estimating a common average treatment effect across all treated regions. In contrast, our empirical strategy enables us to capture not only the average treatment effect of cohesion policy, but also its heterogeneous effects across regions and over time. To this end, we estimate the entire time-varying distribution of regional treatment effects.
This perspective is particularly relevant for evaluating the core objectives of the CF, which are not only to increase the growth of average regional income but, primarily, to reduce economic disparities within the EU. Consequently, the second novelty of this paper is the introduction of standard distributional statistics as a central tool to evaluate the performance of the CF against its policy objectives. Specifically, we employ measures such as the share of regions with positive treatment effect, skewness, kurtosis, average treatment effect by quintile, and the regional Gini coefficient of the distribution over time (e.g., successive seven-year funding cycles).
Building on this framework, we can evaluate the performance of CF in several complementary dimensions with respect to the existing literature (see e.g., Becker et al. (2010), Becker et al. (2013) and Crescenzi and Giua (2020)). To make our results directly comparable to this literature, we start with the average impact on key economic outcomes, such as output and investment. We then analyze the evolution of the distribution of effects to infer convergence dynamics and shifts in regional disparities to identify which types of region benefit most. Finally, we investigate how policy effectiveness varies with transfer intensity, shedding light on potential non-linearities and offering guidance for the optimal design and redistribution of future cohesion interventions.
The empirical literature evaluating the causal effects of EU cohesion policy has so far relied predominantly on Regression Discontinuity Designs or instrumental variable IV strategies. These approaches exploit institutional eligibility thresholds — such as the $75\%$ GDP per capita rule defining access to Objective1~1 or the Convergence Objective or ‘less-developed’ status [^3] Objective 1 or the Convergence Objective or ‘less-developed’ status refers to the classification scheme of the European Commission that determines which regions will receive the lion’s share of the available structural funds. As elaborated in the Council Regulation 2081/93 and the European Commission Decisions 1999/502/EC, 2006/595/EC and 2014/99/EC regions below the threshold of 75% of the average EU GDP per capita belong to this class. — to identify the local causal effects of cohesion transfers on outcomes such as output, investment, or employment (see, e.g., Becker et al. 2010, 2013, 2018, Pellegrini et al. (2013) and Cerqua and Pellegrini (2018)). While such a design provides credible identification, it is typically limited to local average treatment effects for regions close to the eligibility threshold and offers little insight into the dynamic, heterogeneous, or distributional consequences of cohesion spending.
Our focus shifts towards an event-study-type framework, which is more suitable for evaluating dynamic treatment effects and for capturing heterogeneity along several dimensions, e.g., time, regional income levels, transfer intensity and space. However, applying such an approach to the evaluation of cohesion policy raises several empirical challenges. In particular, regional data are characterized by strong spatial and cross-sectional dependencies, while policy impacts are likely to be heterogeneous across regions with different structural characteristics. These complexities must be addressed to obtain credible causal estimates and fully understand the policy’s distributional consequences.
A first empirical challenge arises from spatial dependence across regions, typically driven by geographical proximity and local spillovers. Ignoring these interdependencies can bias treatment effect estimates, as neighboring regions often exhibit correlated outcomes. Several studies have employed standard spatial econometric techniques to address this issue (see, among others, Mohl and Hagen (2010), Bourdin (2018), Crescenzi and Giua (2020), Di Caro and Fratesi (2022) and Amendolagine et al. (2024)), although applications in a causal inference context remain rare, with Crescenzi and Giua (2020) being a notable exception through their spatial RDD approach. A second challenge is cross-sectional dependence, which reflects broader, system-wide sources of correlation across regions. This dependence is often due to unobserved common factors, such as the 2008 Global Financial Crisis or the European sovereign debt crisis, whose effects extend well beyond local neighbourhoods. When these factors are correlated with treatment assignment, they can lead to substantial estimation biases if not properly addressed. This dependence is often due to unobserved common factors, such as the 2008 Global Financial Crisis or the European sovereign debt crisis, whose effects extend well beyond local neighbourhoods. Importantly, these shocks can bias conventional estimators not only when they are correlated with treatment assignment, but also when regions exhibit differential exposure to the same aggregate shock, leading to heterogeneous comovements in outcomes. Both channels generate substantial estimation bias if not properly accommodated.
Finally, regional policies are unlikely to exert homogeneous effects across space. Yet, much of the existing literature implicitly assumes identical treatment effects across regions, overlooking important heterogeneity linked to income levels, structural characteristics, and initial economic conditions. Recognising and modelling such heterogeneity is crucial not only for understanding the full distributional consequences of cohesion policy but also for assessing how its effectiveness evolves over time and across different types of regions. Our approach explicitly addresses these issues by estimating region-specific, time-varying effects, thereby providing new insights into the differential impacts of cohesion spending and informing the design of more targeted and effective policy interventions.
We address these challenges by employing a matrix completion approach implemented within a factor model framework (Xu, 2017, Athey et al., 2021). This approach generalizes both the conventional Two-Way Fixed Effects (TWFE) model and the synthetic control (SC) method (Abadie et al., 2010), thereby combining their main advantages while relaxing key identifying assumptions. In particular, by modelling the data-generating process through interactive fixed effects (Pesaran, 2006, Bai, 2009a), we can explicitly account for unobserved time-varying common factors — potentially correlated with treatment assignment — that would otherwise bias conventional DiD estimates (Gobillon and Magnac, 2016). This feature is especially relevant in our setting, where regional data are relatively scarce and where such confounders, if ignored, would invalidate the parallel trends assumption central to standard DiD approaches (Angrist and Pischke, 2009, Sun and Abraham, 2021), De Chaisemartin and d’Haultfoeuille (2024). Our framework also extends beyond classical synthetic control methods by allowing for staggered treatment adoption across multiple regions and for heterogeneous factor loadings, which capture region-specific responses to common shocks. As a result, we are able to recover region-specific, time-varying treatment effects in the presence of complex cross-sectional dependencies and heterogeneous dynamics — an innovation in the context of cohesion policy evaluation. This richer modelling approach not only strengthens causal identification but also enables a more nuanced assessment of the Cohesion Fund’s effectiveness across time, space, and the income distribution.
Our findings show that, on average, the CF has a sustained positive impact on both GVA and GFCF in recipient regions, relative to the counterfactual where the EU does not provide funds. This is consistent with previous research, see for example, Becker et al. (2010) and Pellegrini et al. (2013) who document a positive impact of the structural funds on growth. The CF effect is front-loaded, reaching its peak within the first seven years after a region is included in the program. For the average ‘hypothetical’ treated EU region, GVA per capita is about €815 higher after seven years and €2961 higher after fourteen years than in the counterfactual scenario. GFCF per capita is €395 and €680 higher over the same periods, respectively.
These findings are further supported by distributional metrics such as the average treatment effect by quintile and what we call the regional Gini coefficient. Quintiles are defined by regional GVA per capita in the year prior to treatment. We estimate that regions in the lowest (highest) quintile exhibit the largest (smallest) average treatment effects. Similarly, the regional Gini coefficient—which captures cross-regional dispersion in GVA per capita—declines from about 31% to 28% over the period 1995–2022. In sharp contrast, under the counterfactual where we assume the absence of CF program, there is no reduction in the dispersion of GVA across regions; in fact, we estimate a modest increase over time. Similar patterns emerge for investment (GFCF), with even stronger and more uneven effects that are highly concentrated among less developed regions.
Finally, we find that the relationship between treatment intensity, measured as the share of CF expenditures to GVA in real terms, and economic impact is clearly non-linear across all time horizons. After the seventh year under treatment, the optimal CF intensity is approximately 0.6% of GVA, at which point the estimated effect reaches 25%. Beyond this threshold, additional funding produces diminishing economic returns. When the horizon is fourteen or twenty one years the optimal intensity is estimated at 0.7% and 0.86%, yielding maximum effects of 0.32% and 0.47%, respectively. These results support the hypothesis of a maximum desirable level of transfers proposed by Becker et al. (2012), suggesting that while moderate levels of EU funding can stimulate growth, when transfers exceed a threshold, they can become inefficient or even counterproductive.
@@ 2 Related literature
The present paper contributes to the empirical literature on the ex-post evaluation of cohesion policy transfers. In recent years, this literature has focused on the application of causal inference techniques. [^4] There exists a large empirical literature that has employed more traditional (spatial) econometric approaches to identify the effects of cohesion policy, focusing either on the long-run effects or on the calculation of multipliers. More details on this strand are available in Appendix A and the bottom panel of Table LABEL:tab:litreview By calculating counterfactual scenarios and comparing them with the actual observed outcomes, they arrive at the net causal impact of the cohesion transfers. A schematic review of the relevant literature can be found in the top panel of Table LABEL:tab:litreview.
Following the seminal contribution of Becker et al. 2010, a number of papers have leveraged on a specific institutional characteristic of the cohesion transfers programme to setup their research design. Specifically, the European Commission classifies regions as attaining the ‘Objective I’ or ‘Convergence Objective’ or ‘less developed’ status if their GDP per capita (in PPS) prior to the commencement of each budget cycle is less than 75% of the EU average. [^5] Indicatively, for the 2014-2020 period the list of regions belonging to the ‘less developed’ group was decided using data for the 2007 - 2009 period, as detailed in the Commission Implementing Decision 2014/99. This classification in turn determines the amount of cohesion transfers each region is eligible for; the ‘less developed’ regions receive the bulk of the transfers. [^6] As such, all regions receive at least some funding; eligibility or not for the ‘less developed’ status determines the size of the transfer. As such, the focus is on whether being assigned the ‘less developed status has a causal impact on economic performance.
The analysis of Becker et al. 2010, 2012, 2018, Gagliardi and Percoco 2017 and Crescenzi and Giua 2020 rests on a fuzzy [^7] The fuzzy design arises due to the fact that some regions that were formally eligible for receiving the ‘less developed’ status were not included in this group of regions and vice versa. RDD setup and, broadly, points to an overall positive causal effect. A positive effect is also identified in Pellegrini et al. (2013) who use a ‘sharp’ RDD approach. Interestingly, all these contributions examine the impact of all the available policy instruments simultaneously [^8] That is, the European Regional Development Fund, the European Social Fund and the Cohesion Fund. and largely do not distinguish between programming periods. Lastly, in an interesting extension, Becker et al. (2013) estimate a heterogeneous local treatment effect based on an extension of the standard RDD that allows for the examination of the impacts of each region’s absorptive capacity.
@@
A separate branch of this literature shifted the focus to the more granular issue of whether the level of transfer intensity that regions receive is what matters for economic performance; that is, whether higher levels of treatment lead to a higher or lower impact on the outcome compared to lower levels of treatment. In this context, Mohl and Hagen (2008), Becker et al. (2012) and Cerqua and Pellegrini (2018) employ a generalized propensity score approach following Hirano and Imbens (2004) and estimate the so-called dose-response function. This allows the examination of whether there exists either a minimum necessary level or a maximum desirable level of regional transfer intensity necessary to achieve positive effects from cohesion policy expenditures. Both Becker et al. (2012) and Cerqua and Pellegrini (2018) identify the presence of non-linear inverted-U shaped relationship between transfers and economic performance.
Our work complements and extends these contributions by adopting a generalized synthetic control framework rooted in the matrix completion and interactive fixed effects literature (Gobillon and Magnac 2016, Athey et al. 2021; Xu 2017, 2023). This approach addresses several well-known limitations of RDD-, IV-, and dose–response-based evaluations. First, whereas RDD and IV designs identify only local or complier-specific effects around the eligibility threshold, our estimator yields region- and time-specific causal effects for the full set of treated units. This enables the analysis of the dynamic evolution of policy impacts over the entire treatment horizon and across the full distribution of regions, rather than a single average homogeneous effect. Second, unlike dose–response models, our estimator does not require unconfoundedness or common support conditions and avoids the high finite-sample variance associated with nonparametric identification. Third, by incorporating interactive fixed effects, our framework explicitly models unobserved time-varying confounders and heterogeneous exposure to common shocks, such as the 2008 Global Financial Crisis or the European sovereign debt crisis; these are factors that can bias conventional RDD/IV or DiD estimators but these approaches cannot adequately accommodate.
Beyond these methodological advantages, our framework enables us to exploit the full 1993–2022 period during which the Cohesion Fund has operated and to obtain annual causal effects rather than effects aggregated over programming periods. This allows us to examine the dynamics of adjustment, assess the extent to which Cohesion Fund transfers mitigated large macroeconomic shocks, and study whether the relationship between transfer intensity and regional outcomes exhibits a non-linear pattern. Finally, by focusing on the Cohesion Fund—the instrument targeted at the EU’s relatively poorer Member States—we shed new light on the heterogeneous effectiveness of cohesion spending across space, time, and the regional income distribution.
We note here that there exists a separate, voluminous empirical literature related to the identification of the impact of cohesion policy. This literature focuses on more traditional (spatial) econometric approaches and is surveyed in Appendix A; for a schematic review see the bottom panel of Table LABEL:tab:litreview. Indicatively, Canova and Pappa (2025) provide evidence on the size of the multipliers of the cohesion policy instruments, noting the heterogeneity arising due to, for example, location and the level development; Amendolagine et al. (2024) employ a heterogeneous panel model and show that the impact of cohesion Policy is positive but highly heterogeneous, while also generating spatial spillover effects.